OpenStudy (anonymous):

Factor completely 3x^2+ 12x+ 7 Help please

OpenStudy (shamil98):

Single out 3x^2 + 12x and find the gcf 3x( x + 4) + 7 that's the simplest form.

hartnn (hartnn):

do you know what the determinant is ? \(b^2-4ac\) heard of it ?

OpenStudy (shamil98):

you mean discriminant? @hartnn

hartnn (hartnn):

oh, lol yes...discriminant :P

OpenStudy (shamil98):

It can't be factored out in (x+ ) (x + ) so the gcf and reduce to simplest form is the best approach i think.

hero (hero):

Using the equation \(\left(\dfrac{b}{2} + d\right)\left(\dfrac{b}{2} - d\right)=ac\) we have: \((6 + d)(6 - d) = 21\) \(36 - d^2 = 21\) \(36 - 21 = d^2\) \(15 = d^2\) \(\pm\sqrt{15} = d\)\ \((6 + \sqrt{15})(6 - \sqrt{15}) = 21\) \((6 + \sqrt{15}) + (6 - \sqrt{15}) = 12\) Now input the last row into the quadratic equation in place of 12: \(3x^2 + ((6 + \sqrt{15}) + (6 - \sqrt{15}))x + 7\) \(3x^2 + (6 + \sqrt{15})x + (6 - \sqrt{15})x + 7\) Factor by grouping: \(3x^2 + (6 + \sqrt{15})x + (6 - \sqrt{15})x + \dfrac{(6 + \sqrt{15})(6 - \sqrt{15})}{3}\) \(3x\left(x + \dfrac{6 + \sqrt{15}}{3}\right) + (6 - \sqrt{15})\left(x + \dfrac{6 + \sqrt{15}}{3}\right)\) \((3x + 6 - \sqrt{15})\left(x + \dfrac{6 + \sqrt{15}}{3}\right)\) Annoying and abstruse, but it does factor.

hartnn (hartnn):

i meant it does not factor into integral terms

hartnn (hartnn):

when b^2-4ac is not a perfect , we cannot factor it as (ax+b)(cx+d) where, a,b,c,d are integers

hero (hero):

It said factor completely so I didn't make any assumptions about factoring limitations. Usually it will say factor if possible.

hartnn (hartnn):

ohh

hero (hero):

That being said, the user probably revised the original question given to him or her.

OpenStudy (mertsj):

It is prime